i have algebra with trigononmetry. my book is called Algebra and trigonometry structure and method, book 2 by Mcdougal Littell.the section is on how to translate word phrases into algebraic expressions and word sentences int equations. the directions say to express each answer in simplest form in terms of the given varialbe.the problem i need help on is this one : " the length of,width,and height of a rectangular box are consecutive intergers, and the largest dimension is k cm. find the volume V of the box. ( hint V= lwh )"
Answer by edjones(8007)

let the dimensions be k, k-1, k-2
V=k(k-1)(k-2)=k(k^2-3k+2)
=k^3-3k^2+2k=V
Ed